ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 144, 2004

Bell, J. P. Dirichlet series whose partial sums of coefficients have regular variation

343

Burgstaller, B. Slightly larger than a graph C*-algebra  1

Deza, M., M. Dutour and M. Shtogrin On simplicial and cubical complexes with short links

109

Dodos, P. Dichotomies of the set of test measures of a Haar-null set  15

Dutour, M. See M. Deza

Gonzalez-Aviles, C. D. Finite modules over non-semisimple group rings

61

Hsu, T. and D. T. Wise Groups with infinitely many types of fixed subgroups

93

Kuhn, M. G. and T. Steger Free group representations from vectorvalued multiplicative functions, I

317

Lindenstrauss, E. and K. Schmidt Invariant sets and measures of nonexpansive group automorphisms

29

Lubotzky, A. and B. Martin Polynomial representation growth and the congruence subgroup problem

293

Martin, B. See A. Lubotzky

Melbourne, I. and A. Török Statistical limit theorems for suspension flows

191

Mukhtarov, O. Sh. and E. Tunç Eigenvalue problems for Sturm– Liouville equations with transmission conditions

367

Narayanan, E. K. and S. Thangavelu An optimal theorem for the spherical maximal operator on the Heisenberg group

211

Novo, S., R. Obaya and A. M. Sanz Almost periodic and almost automorphic dynamics for scalar convex differential equations

157

Obaya, R. See S. Novo

Pakovich, F., N. Roytvarf and Y. Yomdin Cauchy-type integrals of algebraic functions

221

Roytvarf, N. See F. Pakovich

Sanz, A. M. See S. Novo

Schmidt, K. See E. Lindenstrauss Shtogrin, M. See M. Deza Sodin, M. and B. Tsirelson Random complex zeroes, I. Asymptotic

normality

125

Steger, T. See M. G. Kuhn

Subaşi, M. and B. Yildiz A stability estimate for an elliptic problem 151

Thangavelu, S. See E. K. Narayanan Török, A. See I. Melbourne Tsirelson, B. See M. Sodin Tunç, E. See O. Sh. Mukhtarov Wise, D. T. See T. Hsu Yildiz, B. See M. Subaşi Yomdin, Y. See F. Pakovich